OpenStudy (anonymous):

if f(x)=(sqrt(x)) AND g(x)=(sqrt(2-x)) what does (f+g)(x)= So this is the entire problem: f(x)=[sqrt(x)], g(x)=[sqrt(2-x)] 1. find the domain of the following: f, g, f+g, fg, ff, f/g, g/f 2. find: (f+g)(x), (f-g)(x), (fg)(x), (ff)(x), (f/g)(x), (g/f)(x)

OpenStudy (anonymous):

(f+g)(x) = f(x) + g(x)

OpenStudy (anonymous):

what about the f(x)=\[\sqrt{x}\] and g(x)=\[\sqrt{2-x}\] ?